Executive Chalk is financed solely by common stock and has outstanding 42 million shares with a market price of $44 a share. It now announces that it intends to issue $660 million of debt and to use the proceeds to buy back common stock.

  1. a. How is the market price of the stock affected by the announcement?

  2. b. How many shares can the company buy back with the $660 million of new debt that it issues? (Enter your answer in millions.)

  3. c-1. What is the market value of the firm (equity plus debt) after the change in capital structure? (Enter your answer in millions.)

  4. c-2. Did the market value of the firm change?

  5. d. What is the debt ratio after the change in structure? (Round your answer to 2 decimal places.)

  6. e. Who (if anyone) gains or loses?
Answer #1

(a): Market price of the stock will not be affected by the announcement.

(b): Total amount of debt = $660 million. Market value of share = $44. No. of shares that can be bought back = $660 million/$44

= 15 million shares

(c -1): Total no. of shares outstanding = 42 million. No. of shares that can be bought back = 15 million. No. of shares outstanding after buyback = 42-15 = 27 million.

Market value = debt + equity. Debt = $660 million. Equity = 27 million * $44 = 1188. Thus value = 660+1188

= $1,848 million

(c-2): No there is no change in market value of the firm.

(d): Debt = 660 and equity = 1188. Thus debt ratio = 660/(660+1188) = 0.36

(e): No one will gain or lose anything.